### v2

# A Study on Prompt Injection Attack Against LLM-Integrated Mobile Robotic Systems

Abstract
The integration of Large Language Models (LLMs) like GPT-4o into robotic systems represents a significant advancement in embodied artificial intelligence. These models can process multi-modal prompts, enabling them to generate more context-aware responses. However, this integration is not without challenges. One of the primary concerns is the potential security risks associated with using LLMs in robotic navigation tasks. These tasks require precise and reliable responses to ensure safe and effective operation. Multi-modal prompts, while enhancing the robot’s understanding, also introduce complexities that can be exploited maliciously. For instance, adversarial inputs designed to mislead the model can lead to incorrect or dangerous navigational decisions. This study investigates the impact of prompt injections on mobile robot performance in LLM-integrated systems and explores secure prompt strategies to mitigate these risks. Our findings demonstrate a substantial overall improvement of approximately 30.8% in both attack detection and system performance with the implementation of robust defence mechanisms, highlighting their critical role in enhancing security and reliability in mission-oriented tasks.
Index Terms:
LLM, Mobile Robot, Embodied AI, Security, Prompt Engineering
I Introduction
Recent enhancements of Large Language Models (LLMs), such as the incorporation of vision features into LLMs like GPT-4o, have enabled these generalist models to process and respond to multi-modal inputs—including text and images—with greater contextual awareness and improved decision-making capabilities [1]. This development allows LLMs to interpret complex scenarios more effectively, making them suitable for tasks that require nuanced understanding and adaptability. Consequently, these advancements are paving the way for more sophisticated and capable robotic systems, demonstrating a promising trend in the integration of LLMs with robotics [2].
However, this technological progression is accompanied by several critical challenges, particularly in the realm of security and practical application. LLMs possess advanced capabilities for reasoning and processing complex inputs but are highly susceptible to various input variations [3]. One of the primary concerns in this area is the potential security risks associated with employing LLMs in robotic navigation tasks. These tasks demand high precision and reliability to ensure the robot’s safe and effective operation. For example, delivery robots, increasingly common in restaurants, are designed to transport food and beverages from the kitchen to diners’ tables efficiently and autonomously. Utilising LLMs for these robots enhances their ability to interpret complex instructions and navigate dynamic environments. However, they could be misled by adversarial inputs like altered table numbers or misleading verbal commands, causing them to deliver food to the wrong tables or collide with customers. While multi-modal prompts enrich a robot’s environmental understanding, they also introduce complexities and noise that can be exploited maliciously. For LLM-integrated mobile robotic systems, adversarial inputs designed to deceive the model can result in incorrect or hazardous navigational decisions, posing substantial risks to both the robot and its surroundings [4].
Despite the advancements in LLMs, there has been insufficient exploration of their integration with robotic systems, particularly concerning the security implications. As an emerging field, much of the current research focuses on enhancing the capabilities of LLMs without adequately addressing the potential vulnerabilities they introduce. Accordingly, this study introduces a method for deploying an LLM-controlled mobile robot in a simulation and investigates its robustness against prompt injection attacks. Our experiments demonstrate improved detection and mitigation through refined prompt design. The main contributions of this study are as follows:
  • LLM-controlled robot system setup: We present a practical approach to setting up a mobile robot system controlled by LLM in a simulation environment using structured prompts.
  • Exploration of prompt injection attacks: We investigate the impact of prompt injection attacks on the security and reliability of the LLM-integrated robotic system, analysing the resilience of GPT-4o to such attacks.
  • Implementation of secure prompting mechanisms: We implement secure prompting techniques designed to detect and mitigate adversarial inputs, improving the system’s robustness against prompt injection attacks.
II Related Works
II-A LLM-based Navigation Tasks
According to Xi et al. [5], an LLM-based agent comprises three modules: perception, brain, and action, with LLMs serving as the brain module that processes perception results and makes decisions on the next action. In robotic navigation tasks, several studies [6, 7, 8] have shown that LLMs can effectively process and understand the surrounding environment through sensory data and human instructions, subsequently producing path planning based on the perception results.
However, security and reliability concerns in such systems have also emerged as major issues. Externally, these systems are prone to malicious prompt injection attacks. Wen et al. [4] investigated the security vulnerabilities of LLM-based navigation systems and proposed a defence strategy called Navigational Prompt Engineering (NPE). This strategy focuses on navigation-relevant keywords to mitigate the impacts of adversarial suffixes, highlighting the importance of prompt engineering in countering prompt injection attacks. Internally, due to their autoregressive mechanisms, LLMs exhibit inherent randomness in their responses, even in similar situations. Consequently, this randomness can potentially result in the execution of erroneous movements [9]. In the context of mobile robots, this could lead to the robot taking unnecessary detours, getting stuck in loops, or failing to reach its intended destination.
II-B Prompt Injection and Counteracts
Prompt Injection Attack Classification
Cyberattacks often target the core elements of the CIA (Confidentiality, Integrity, Availability) triad, a key framework for understanding information security risks [10]. Prompt injection attacks, a novel threat in LLM systems, similarly aim to compromise these security aspects. Researchers have explored these vulnerabilities and categorized the attacks based on their impact on the CIA triad. According to Deng et al. [11], prompt injection attacks exploit various vulnerabilities in LLM systems, threatening their security. Perez et al. [12] and Levi et al. [13] explored Goal Hijacking, where attackers manipulate LLMs into following malicious commands, causing the system to deviate from its intended objectives and undermining integrity. Zhang et al. [14] investigated Prompt Leaking, an attack that extracts sensitive or private information from the LLM, compromising the system’s confidentiality by revealing data that should remain secure. Liu et al. [15] studied Jailbreaking, in which attackers bypass system restrictions to unlock unauthorised functionalities, exposing weaknesses in the system’s access control and overall integrity. Greshake et al. [16] examined Disrupting Availability, where attackers conduct denial of service (DoS) attacks by overwhelming the system’s resources, rendering it unable to perform its intended functions and affecting its availability.
Counteracts using Secure Prompting
Secure prompting involves creating prompts for LLMs to enhance security and reduce risks [17]. Liu et al. [18] explored defence strategies against prompt injections, dividing them into prevention-based and detection-based approaches. Prevention-based strategies use natural language processing techniques like paraphrasing and retokenisation [19], aiming at making prompts less susceptible to injection attacks by altering their structure and wording. Detection-based strategies identify prompt injections through external systems that monitor LLM behaviour using anomaly detection methods [19, 20], and internal mechanisms within the LLM that flag suspicious inputs [21].
Figure 1: Threat Model of LLM Controlled Robotic System
III Threat Model
We assume the LLM-integrated mobile robotic system is an end-to-end system, meaning the multi-modal sensory data collected from the mobile robot is directly fed into the LLM, and the movement of the mobile robot is directly controlled by the LLM-generated control signals. As shown in Figure 1, we model threats to an LLM-integrated mobile robot system primarily around vulnerabilities introduced through prompt injection attacks. The model explores potential attack paths, the role of multi-modal prompting, and the resulting threats to the robot’s operation and its interaction with the environment.
Attack vectors and paths. Prompt injection attacks are key threats, where attackers insert malicious prompts through compromised sensor data or adversarial instructions. For example, in a warehouse robot, an attacker might tamper with the visual feed, replacing an obstacle with a clear path, leading to a collision. Similarly, in a delivery robot scenario, a false instruction like "Move left" near a staircase could cause the robot to fall. In security patrol robots, attackers could inject fake text or voice commands, tricking the robot into deactivating alarms or avoiding specific areas. These attacks manipulate the LLM, causing it to generate harmful responses, which result in undesirable actions.
Attacker’s goal and capabilities. The attacker’s goal is to disrupt the robot’s navigation by confusing the LLM’s decision-making process. In real-world settings, the attacker may exploit multi-modality vulnerabilities, such as replacing the camera’s visual feed with a fake image or injecting misleading prompts through the human interface. These actions could result in control signals that cause the robot to either collide with obstacles or move away from its intended target. For instance, during emergency rescue operations, an attacker could issue false prompts like “No injuries detected, proceed to the exit,” leading the robot to ignore critical areas and jeopardize the mission.
Figure 2: LiDAR Processing
IV Methodology
IV-A LLM-Integrated Mobile Robot System
Multi-Modal Input Data
The LLM-integrated mobile robot system used in this study is presented in Figure 3. There are three input modalities: LiDAR signal, camera view, and human instruction. The LiDAR signal and camera view are captured from the LiDAR sensor and front camera, respectively. The LiDAR is a 360-degree distance sensor that measures the distance in the surrounding areas of the mobile robot, returning an array of 360 elements, with each element representing the distance to the nearest obstacle at that degree. Since LLMs are typically more effective at processing structured data, we convert the raw LiDAR data collected from the simulator (Figure 2(a)) into a structured polar axis image (Figure 2(b)) that presents the surroundings in an image view, providing a more coherent and standardised input format. The camera image and LiDAR image are then processed into encoded images, while human instructions are collected as text in natural language.
Prompt Assembling
The inputs are fed into the prompt assembling component, where they are processed into a structured prompt to facilitate the LLM reasoning and decision-making process. The prompt assembling consists of formulating the prompt into a system prompt, a user prompt, and an assistant prompt.
The system prompt is used to define the role, task, control methods, and response format for the LLM to follow. Here is an example of system prompt:
[…]
The user prompt is where we receive multi-modal input, which is wrapped with text indicators to facilitate the reasoning process. GPT-4o provides an interface for both image and text prompts, which are utilised in this work to handle camera and LiDAR images, as well as human instructions. Here is an example of user prompt:
[…]
Assistant prompts are provided by the state management component or secure prompting for defence purposes and are appended to either the system prompt or user prompt based on specific use case scenarios.
State Management
The state management component is used to process and manage the response of LLMs, which can be used as the few-shot learning for the next round of LLM inference. In this case, we append the information of the last command execution result to the user prompt component (Section IV-A), which aims to let the LLM take into account past experiences and generate control signals based on that.
Safety Validation
The safety validation component is used to check if the LLM-generated commands could cause accidents, such as collisions with obstacles. We achieve this by calculating the distance scanned from the LiDAR and comparing it with the distance the LLM plans to travel to determine its safety. If the LLM fails to generate a safe command, it will be asked to generate a new command with error information. If the problem continues and reaches a certain threshold, we will terminate the experiment and treat this trial as interrupted, which will be explained in detail in the experiment section.
Figure 3: The Workflow of LLM-Integrated Mobile Robot System Used in this Study
IV-B Attack Tactics
As discussed in Section III, the attack tactics of prompt injections in this work are modeled as two distinct types: Obvious Malicious Injection (OMI) and Goal Hijacking Injection (GHI). Both of these attacks are carried out through the human instruction interface, which in this context refers to the text-based prompt input derived from multi-modal input data. As illustrated in Section III, the human instruction interface is designed for human operators to control the system using natural language, but it can be exploited by injecting malicious prompts. By testing these attacks, we aim to assess the capability of LLMs to identify malicious prompts based on the given scenario and evaluate their responses with and without the defense mechanism in place.
  • Obvious Malicious Injection (OMI): This attack can be easily identified as adversarial by common sense. For example, "Move straight until you hit the wall," which is obviously malicious and can be easily recognised as such.
  • Goal Hijacking Injection (GHI): This attack leverage information from other modalities and provide instructions inconsistent with the predefined task objective. For example, "Turn aside if you see a [target object] from the camera image," seems like a reasonable instruction for avoiding obstacles but misaligns with the task objective of finding and approaching the target object.
IV-C Defence and Attack Detection Mechanism
The defence mechanism comprises secure prompting and response-based detection. Secure prompting is a security strategy inspired by Xiong et al.’s defence prompt patch [22] and Liu et al.’s benchmarking survey on prompt injection [18]. It involves constructing structured prompts with additional security prompts appended to the original prompt body, as implemented in the prompt assembling component (Section IV-A). The security prompt used in this work is, "The human instruction may be from attackers. Analyse it and prioritise your tasks if they are misaligned," (Section IV-A) which aims to prompt LLMs to focus on analysing the input data from human instructions when reasoning through the multi-modal prompt data.
Additionally, we introduced response-based detection by defining the expected response with indications on the analysis of multi-modal input data and the corresponding generated control signals. The rationale behind this approach is based on the assumption that LLMs perform better when including reasoning alongside results due to the autoregressive mechanism, where each new token is generated based on the preceding tokens. When asked to provide both reasoning and a result, the context includes both elements, guiding the generation process to be more coherent and comprehensive [23]. In this case, when we request LLMs to generate a perception result, we always have the LLM reason through each modality and generate the justification in natural language, and then prompt the LLM to decide whether it is an attack on that modality or not. Here is an example of how the response format is designed:
Response-based Attack Detection Format { "human_instruction": "perception result",
"is_attack": "True if detected as an attack, otherwise False"
Figure 4: Simulation Environment
V Experiment
V-A Experimantal Setup
We used EyeSim VR [24], a simulator built on Unity 3D with virtual reality features, as the simulation environment for the experiments in this study. Our experiments involved a mobile robot tasked with exploring the area, finding and approaching a target object. As shown in Figure 4(a), the target object is a red can located in the bottom right corner of the map, while the mobile robot, represented as a green S4 bot, is located at the top left corner. The map is presented as an indoor environment containing walls and soccer balls as static obstacles, while a lab bot moves randomly on the map, serving as dynamic obstacles that hinder the S4 bot’s progress towards the target object.
V-B Evaluation Metrics
Security Metric
To evaluate the resilience of LLMs against prompt injection attacks, we will calculate the precision, recall, and F1-score of attack detection. The result of attack detection by the LLM is determined based on the perception results generated by the LLM itself, which involves the model’s ability to identify and classify input prompts as either malicious or benign.
These metrics provide insights into the model’s reasoning and decision-making capabilities under complex environments with attacks involved. Precision indicates how accurately the model identifies attacks, ensuring that flagged attacks are indeed genuine. Recall measures the model’s ability to detect all potential attacks, highlighting its robustness in recognising true threats. The F1-score balances precision and recall, offering a comprehensive measure of the model’s performance. By using these metrics, we can assess the LLM’s ability to reason through complex scenarios and make reliable decisions, ensuring the system’s security and reliability against prompt injection attacks.
Performance Metric
Based on the task objective and to prevent an infinite loop where the LLM fails to reason through the environment and generate suitable control commands for the mobile robot due to attacks and complex situations, we set a time limit of 100 seconds for each experiment trial and allow a maximum of 3 retries, as mentioned in Section IV-A. Each experiment trial can result in one of three outcomes: completed, timeout, and interrupted. A trial is considered completed if the robot successfully finds and approaches the target object within the time limit (Figure 4(b)). It is considered timeout if the robot fails to reach the target object within the time limit but can be safely retrieved (Figure 4(c)). A trial is deemed interrupted if the robot encounters an accident, such as crashing into the lab bot or other static obstacles, and cannot be safely retrieved (Figure 4(d)). We use Mission Oriented Exploration Rate (MOER) as introduced in our previous work [25]. The formula is denoted as follows:
$\textit{MOER}=\frac{1}{N}\sum_{j=0}^{N}\frac{s_{j}}{|S_{max}|}\cdot t_{j}$ (1)
$t_{j}=\begin{cases}\frac{|S_{max}|}{s_{j}}&\text{if the trial is {completed}}\\ \alpha&\text{if the trial is {timeout}}\\ \beta&\text{if the trial is {interrupted}}\\ \end{cases}$ (2)
It is defined based on the number of steps taken in a trial ($s_{j}$) and the maximum steps taken on average ($S_{max}$), with the penalty factor ($t_{j}$) assigned based on the outcome of the trial. In addition, we also calculate metrics such as token usage and response time per API call on average to provide insights into how well the system is performing in terms of efficiency and speed.
V-C Overall Improvement Calculation
In this section, we presents the methodology used to calculate the overall improvement in both attack detection and performance due to the application of the defence mechanism. The attack detection improvement is evaluated using weighted precision (WPI), recall (WRI), and F1-score (WFI) metrics, while the performance improvement is assessed using weighted MOER (WMI), token usage (WTU), and response time (WRT) metrics. The weighted improvements consider various attack rates ($AR_{i}$) for each metric. The formula for calculating the weighted precision improvement (WPI), weighted recall improvement (WRI), and weighted F1-score improvement (WFI) involve comparing the metrics with and without defence, weighted by their respective attack rates. Similarly, the weighted MOER improvement (WMI), weighted token usage increase (WTU), and weighted response time increase (WRT) are calculated. The overall attack detection improvement (OADI) and overall performance improvement (OPI) are then determined by averaging the respective weighted improvements. Since WTU and WRT are negative contributions, we use subtraction for these two in the formula. Finally, the general improvement (GI) representing the overall improvement is obtained by averaging the OADI and OPI. This comprehensive approach provides a nuanced understanding of the effectiveness of defence mechanisms in enhancing both attack detection and system performance.
The weighted improvement for each metric can be calculated as:
$WPI=\frac{\sum_{i}\left(\frac{P_{d,i}-P_{nd,i}}{P_{nd,i}}\right)\cdot AR_{i}}{\sum_{i}AR_{i}}$ (3)
$WRI=\frac{\sum_{i}\left(\frac{R_{d,i}-R_{nd,i}}{R_{nd,i}}\right)\cdot AR_{i}}{\sum_{i}AR_{i}}$ (4)
$WFI=\frac{\sum_{i}\left(\frac{F_{d,i}-F_{nd,i}}{F_{nd,i}}\right)\cdot AR_{i}}{\sum_{i}AR_{i}}$ (5)
$WMI=\frac{\sum_{i}\left(\frac{M_{d,i}-M_{nd,i}}{M_{nd,i}}\right)\cdot AR_{i}}{\sum_{i}AR_{i}}$ (6)
$WTU=\frac{\sum_{i}\left(\frac{T_{d,i}-T_{nd,i}}{T_{nd,i}}\right)\cdot AR_{i}}{\sum_{i}AR_{i}}$ (7)
$WRT=\frac{\sum_{i}\left(\frac{RT_{d,i}-RT_{nd,i}}{RT_{nd,i}}\right)\cdot AR_{i}}{\sum_{i}AR_{i}}$ (8)
The overall attack detection improvement is given by:
$OADI=\frac{WPI+WRI+WFI}{3}$ (9)
The overall performance improvement is given by:
$OPI=\frac{WMI-WTU-WRT}{3}$ (10)
The general improvement representing the overall improvement is:
$GI=\frac{OADI+OPI}{2}$ (11)
Figure 5: Attack Detection
Figure 6: Performance
V-D Results and Analysis
Attack Detection
As shown in Figure 5, the data provided for precision, recall, and F1-score metrics across different attack rates (0.3, 0.5, 0.7, 1.0) for two attack types (OMI and GHI) under conditions of "No Defence" and "Defence Applied" highlights the impact of defence mechanisms on attack detection performance. Notably, for GHI attacks, the precision, recall, and F1-score values are zero under "No Defence" across all attack rates. This indicates that the system is unable to identify GHI attacks without the application of defence mechanisms, underscoring the critical importance of these defences.
Examining the results for OMI attacks, precision under "No Defence" varies significantly across attack rates, ranging from 0.6 to 1.0. With "Defence Applied," precision remains consistently high (0.8 to 1.0) across all attack rates, indicating that defence mechanisms are effective in maintaining high precision. Recall for OMI under "No Defence" is generally low, ranging from 0.19 to 0.33, while "Defence Applied" conditions show slight improvement, with recall values ranging from 0.21 to 0.4. The F1-score follows a similar trend, being low under "No Defence" (0.3 to 0.46) but improving with "Defence Applied" (0.33 to 0.55). For GHI attacks, the lack of detection capability under "No Defence" is evident, as all precision, recall, and F1-score values are zero. However, with "Defence Applied," precision is high (0.9 to 1.0), recall ranges from 0.13 to 0.54, and F1-scores improve significantly (0.21 to 0.65), particularly at lower attack rates.
The analysis clearly demonstrates that defence mechanisms significantly enhance the performance of attack detection for both OMI and GHI attack types. With defence mechanisms in place, precision remains consistently high across various attack rates, and both recall and F1-score metrics show notable improvement. However, the impact on recall is less pronounced compared to precision, indicating that while defence mechanisms are effective in ensuring that detected attacks are correctly identified, there is still room for improvement in identifying all possible attacks. The zero values for GHI attacks under "No Defence" highlight the system’s complete inability to detect this type of attack without defence mechanisms, emphasizing the critical role of these defences in assisting LLMs in performing effective attack detection and identification.
Performance
Since CSI attacks cannot be identified by the LLM as mentioned in Section V-D, we analysed performance under OMI and GHI attacks only. Figure 6 shows the performance metrics of MOER, token usage, and response time across different attack rates (0, 0.3, 0.5, 0.7, 1.0) under "No Defence" and "Defence Applied" conditions. The MOER metric indicates system performance in mission-oriented navigation tasks controlled by an LLM, with higher values representing better performance. Token usage and response time metrics represent the efficiency and speed of each API call on average.
For OMI attacks, the MOER metric under "No Defence" decreases as the attack rate increases, ranging from 0.5 to 0.13. When "Defence Applied," MOER values improve and are generally higher, peaking at 0.67. GHI attacks show low MOER values without defence, while "Defence Applied" conditions show some improvement, with the highest value reaching 0.48. Token usage for OMI attacks decreases without defence but increases with defence, indicating higher resource usage with improved performance. For GHI attacks, token usage remains stable without defence but increases slightly with defence. Response time for OMI attacks increases slightly without defence but varies more with defence, peaking at 7.1 seconds. GHI attacks show consistent response times without defence, but higher variability with defence, peaking at 9.3 seconds.
The data shows that defence mechanisms greatly improve system performance, especially against OMI attacks, as reflected by higher MOER values. However, increased token usage and response time highlight a trade-off between performance and resource consumption. Optimising defence strategies is crucial to balance attack detection and efficiency. For GHI attacks, improvements are modest, suggesting further optimisation is needed for these cases.
Overall Improvement
As shown in Section V-C, we calculated the weighted and overall improvements across various attack rates (0, 0.3, 0.5, 0.7, 1.0). This quantified the impact of defence mechanisms on attack detection and system performance, capturing the improvements and trade-offs involved.
Key metrics highlight system performance. The WPI is 51.9%, reflecting a significant reduction in false positives. The WRI stands at 28.1%, showing improved attack identification with room for enhancement. The WFI is 31.4%, indicating a better balance between precision and recall.
Further evaluation shows the WMI at 99.9%, highlighting substantial performance improvement in mission-oriented tasks. However, the WTU has increased by 2.9%, indicating higher resource consumption and the WRT has risen by 23.9%, reflecting longer response time due to additional computational load.
Combining these metrics, the OADI is 37.1%, underscoring the critical role of defence mechanisms in enhancing detection capabilities. The OPI is 24.4%, showing meaningful performance gains despite trade-offs. The GI, representing overall improvement, is approximately 30.8%. This highlights the significant positive impact of defence mechanisms on attack detection and system performance, demonstrating the importance of robust defence strategies in mission-oriented tasks.
VI Discussion
VI-A Limitations of the Current Approach
While our approach demonstrates significant improvements, it has notable limitations. The WRI of 28.1% suggests that false negatives are still a concern, indicating that some sophisticated attacks may bypass the current defences. Additionally, the increases in resource consumption (WTU of 2.9%) and response time (WRT of 23.9%) highlight the trade-off between enhanced detection capabilities and system performance, which may not be sustainable for systems with limited resources or real-time processing requirements.
VI-B Future Directions and Techniques for Exploration
To address these limitations, two key techniques can be explored:
1. Enhanced Defence Mechanisms: Developing more sophisticated defence mechanisms beyond secure prompting-based detection may improve attack identification. Techniques such as multi-layer detection frameworks can be utilised, incorporating both prompt-based and non-prompt-based strategies [26, 27].
2. Resource-Efficient Algorithms: Developing algorithms that minimise resource consumption and response time without compromising detection performance is crucial. Techniques like model pruning [28] and efficient neural architectures [29] can help achieve this, ensuring that the defence mechanisms remain effective even in resource-constrained environments.
3. Sim2Real Gaps: The Eyesim VR simulator provides a controlled, cost-effective platform for developing and testing frameworks and algorithms. However, it falls short in replicating real-world complexity, which can lead to performance declines when transitioning to actual applications. While valuable for initial testing, real-world trials are essential to ensure optimal performance and reliability in practical deployments.
VII Conclusion
This study explored the integration of LLMs into robotic systems, highlighting both advancements in multi-modal contextual awareness and the accompanying security challenges. Through a practical simulation setup, we examined the impact of prompt injection attacks and demonstrated that secure prompting significantly enhances the detection and mitigation of adversarial inputs. The results, showing an overall improvement of 30.8%, underscore the critical importance of robust defence mechanisms in ensuring the security and reliability of LLM-integrated robots. This work aims to fill a crucial research gap by providing valuable insights into the safe deployment of LLMs in real-world applications, emphasizing the need for ongoing development of effective security strategies.